### v1

# MedicalNarratives: Connecting Medical Vision and Language with Localized Narratives

Abstract
We propose MedicalNarratives, a dataset curated from medical pedagogical videos similar in nature to data collected in Think-Aloud studies and inspired by Localized Narratives, which collects grounded image-text data by curating instructors’ speech and mouse cursor movements synchronized in time. MedicalNarratives enables pretraining of both semantic and dense objectives, alleviating the need to train medical semantic and dense tasks disparately due to the lack of reasonably sized datasets. Our dataset contains 4.7M image-text pairs from videos and articles, with 1M samples containing dense annotations in the form of traces and bounding boxes. To evaluate the utility of MedicalNarratives, we train GenMedCLIP based on the CLIP architecture using our dataset spanning 12 medical domains and demonstrate that it outperforms previous state-of-the-art models on a newly constructed medical imaging benchmark that comprehensively evaluates performance across all modalities. Data, demo, code and models available at https://medical-narratives.github.io/
1 Introduction
Figure 1: MedicalNarratives: Here we show 4 samples from the dataset, these samples come from varying videos containing 4 medical modalities, X-ray, CT, MRI, and Histopathology images from left to right. These samples are selected from interleaved samples extracted from source videos, with each sample showing the representative image captured, the raw input text grounded and aligned in-time with the spatial traces & bbox, the denoised medical and ROI text describing the medical image removing all transcription errors, and non-medical information. Please see more examples across all 4 modalities we capture MedicalNarratives for in the Appendix.
Analyzing medical images is an intrinsically perceptual and conceptual task [86] requiring the expert to extract visual clues from the imaging modalities and to conceptually retrieve knowledge from memory and utilize his/her understanding to lead to a diagnosis. This cognitive process requires connecting spatial regions of the image to clinical understanding, often utilizing a segmental approach to avoid errors. This makes reading and communicating medical imaging information, and innately dense and semantic task requiring understanding and grounding.
Within medical image analysis today, studies on semantic tasks like classification, captioning, and retrieval are typically explored separately from dense tasks like segmentation, detection, and disease monitoring, this can be attributed to the lack of the model architecture that enables exploring both fields of research in tandem, and the lack of datasets to train such models natively. This problem is not unique to medical image analysis; in the general domain this dichotomy exists, even with the similar underlying Transformer architecture [126] model architectures like CLIP [102, 57] perform well on semantic tasks and models like DINO [21] outperform on dense tasks. Recently, there has been active research into developing methods that can perform both types of tasks [138, 76, 42, 100, 84], supported by datasets that contain images, spatial annotations, and the language to connect them [71, 100].
While the cost to collect these datasets can be steep, certain annotation protocols strike a good balance between ease of collection, and utility. Specifically, Localized Narratives [100, 128], proposed a dataset of image, text, and grounding traces. They did this by having annotators describe an image vocally while simultaneously moving a computer mouse to the regions they describe, resulting in holistic descriptions of the scene/image all grounded in real-time with regions of the image, capturing both semantic understanding of the scene and object relations. This protocol of collecting grounded vision-language (VL) datasets does not have strict spatial annotations yet captures strong spatial correlations to the description with every trace, making the protocol uniquely easier to undertake and capture data en mass as it appeals to the human nature to point while describing a scene [61, 129, 46]. Localized narratives have been used to train models on semantic tasks [100, 128, 139], dense tasks [42, 39, 33] and more recently to train generative multimodal language models (MLLM) like PixelLM [134], Quilt-LLaVA [118], and Molmo[32] and other generative image models [69, 136].
Unlike the general domain with web-scale datasets [117] and performative models that translate into products, medical foundational and generative models do not translate well into clinical use. While a major reason is the heightened level of sensitivity and slow adoption, we argue the lack of naturally aligned multimodal data that scales with model parameterization is also a major contributing factor. Recent work like MedTrinity-25M [133], attempts to address this by releasing a multimodal dataset with spatial annotations, but relies on sub-optimally pretrained models to generate text descriptions and Regions of Interests (ROIs) for medical images lacking ground truth annotations, potentially propagating model biases and errors.
To address these limitations, we draw inspiration from how medical experts naturally communicate and teach. In the joint field of cognitive psychology and medical imaging, studying how medical experts analyze patient data, studies leverage the think-aloud protocol [35] to capture data for various types of analysis. Within this protocol, experts verbalize their thoughts as they perform a task, and in certain studies they capture their eye gaze/cursor localizing the regions of the image on which they focus [75, 49]. This protocol has been used to collect some medical datasets [96, 85] including the Tufts dental x-ray database [96] which captures a multimodal dataset including the radiologist’s expertise captured in the form of eye-tracking and think-aloud protocol.
Following this intuition and building upon prior work Quilt-LLaVA [118], we propose MedicalNarratives a dataset that leverages pedagogical medical videos where instructors naturally narrate descriptions while pointing to relevant regions with their cursor - closely mimicking the think-aloud protocol used in clinical practice. Our dataset contains 4.2M image-text pairs across 11 medical modalities and 1 pseudo-medical domain, with interleaving samples between varying modalities (e.g., X-ray and CT for the same patient). Importantly, 875K of these samples are grounded in traces that can be reformatted into bounding boxes or masks, serving to train various tasks including dense models like MedSAM [81, 24] while being flexible for extension in clinical settings with minor adjustments to current diagnostic workflows.
To test the utility of our models we train a vision-language model called GenMedCLIP on our dataset and evaluate it on a new benchmark of datasets that cut across all 11 medical modalities for both classification and retrieval where we see GenMedCLIP outperform prior SOTA models like BiomedCLIP in both tasks with 3% and 14% respectively. We hope future works leverage the dataset to train more grounded generative models similar to Quilt-LLaVA [118], LLaVA-Med++ [133], and PixelLLM [134] as well as open-vocabulary segmentation/detection models. To bolster other use cases we also release the constituting video clips and many other metadata including UMLS entities.
2 Related work
Figure 2: The data curation pipeline for the Video subset of the MedicalNarratives dataset. Search: YouTube video-first search strategy, with filtering by pre-trained classifiers and heuristics. Image: Extracting keyframes of a video, denoising, and identifying medical images. Text: ASR transcription, text correction with LLMs, and medical/ROI text extraction. Traces: Identifying stable chunks of a video, then localizing cursor traces within each chunk. Alignment: Mapping medical/ROI text, traces, and images together. Samples are classified into finer-grained subdomains and samples with discussions of multiple domains are identified with LLMs.
2.1 Vision Language representation
Vision-language (multi-modal) models have evolved over time in both supervised and self-supervised paradigms; in recent studies, contrastive self-supervision objectives [102, 141, 57] that learn by matching paired-modality embeddings have outperformed prior work [79, 74, 26] in downstream tasks and, more importantly, perform better at zeros-shot tasks or on emergent domains for which disparate modalities share a paired domain [40, 145]. In medical imaging, early studies in radiology [141, 51] were pre-trained on specific x-ray images and their reports, and more recently domain specific VL models have pushed the SOTA on various tasks with models developed for Ophthalmology [119], Histopathology [54, 52], Computed Tomography [48], Mammography [25], Dermatology [68], Ultrasound/Echocardiography [27]. These models work well for the specific domains they are trained on and not for other domains, which may not have enough data to train for, hence the push for more general medical models [140, 144, 124].
2.2 Medical (Localized) Narratives
In training these VL models, much research effort is used in sourcing, filtering, and curating medical image(s)-text(s) paired data for pre-training, mostly sourcing general and specific medical domain datasets from Medical reports [60], PubMed [36, 109, 37, 140], books [37], social-media [52, 54], YouTube/videos [54, 118] or mixtures of these [133]. The utilization of these data for dense tasks like segmentation and detection (open-closed vocabulary) is limited as they do not provide any spatial annotation localizing regions of the images to specific labels/text, In contrast, every word in a localized narrative [100, 128, 139, 118] is grounded to a region of the representative image by the point/trace captured. This datasets have been used to train models for semantic reasoning [100, 128, 139], and for dense tasks [42, 39, 33], and they also support training both generative multimodal language models [134, 118, 128] and generative image models [69, 136]. Specifically in medical image analysis, Quilt-LLaVA [118] adopts this paired data structure for training its histopathology chatbots with improved spatial reasoning, and PathNarrative’s [139] hierarchical decision-to-reason localized narrative structure, enables classification and captioning tasks, while providing explainable clues for clinicians to understand and diagnose, improving human-AI collaboration in diagnosis.
3 MedicalNarratives: Curation, Formatting, and Characterization
3.1 Overview
Curating a vision-language dataset with spatial traces from unstrictured pedagogy videos is a non-trivial, as many videos either lack voiced audio, are not in English, fail to contain medically relevant content, or have insufficient medical relevance—for example, present static images of medical content on a slide deck, or briefly cover medical images in pursuit of a different objective. In addition, conventional automatic speech recognition (ASR) systems also struggle with the specialized requirements of medical language transcription, necessitating a non-trivial solution. The de-noising of text and image modalities adds further complexity as the videos are typically conversational and, therefore, inherently noisy. Instructors often record both relevant and irrelevant visual content in their videos, making extracting frames at static intervals non-representative of the medical data contained in the video. To collect MedicalNarratives, we trained models and handcrafted algorithms that leverage the nuances in the instructors’ textual and visual behavior, ensuring accurate collection and alignment of both modalities.
3.2 Sourcing and Filtering Medical Data from PubMed and Videos
The MedicalNarratives curation pipeline expands on methods outlined in Quilt-1M [54] and Quilt-LLaVA [118]. It involves (a) sourcing video/article data across 12 medical imaging domains, (b) filtering videos/articles, (c) denoising the captured images, captions, and trace modalities, and (d) aligning all modalities.
3.2.1 Collecting representative videos and articles.
Videos: We first search for relevant channels and videos on YouTube, focusing on the 12 medical imaging domains including: Magnetic Resonance Imaging (MRI), X-Rays, Computed Tomography (CT), Ultrasound (US), Mammography (Mammo), Surgery (Surg), Endoscopy (Endo), Dentistry, Dermatology, Ophthalmology (Optha), Histopathology (Histo), and General Medical Illustrations. Using keywords collected for each outlined domain (see section 7.1 in the Appendix), we search for relevant videos to expedite discovery and exploit found videos by then searching through the corresponding channel for more relevant videos. Once identified, we download low-resolution versions of the videos for filtering.
Articles: We download article documents from the PubMed Central Open Access Subset (PMC-OA) [90], which contains 5.47M articles. We then process each article, extracting figures and captions. We also search for inline references to figures within the body of an article to add to each figure-caption data pair.
3.2.2 Filtering for narrative medical videos and relevant article samples.
Videos: We modify filtering heuristics set by Quilt-1M to translate to other medical domain videos. These modified heuristics include: excluding videos shorter than $1$ minute in duration or longer than $2$ hours, non-voiced videos, and videos from channels with more than 1M subscribers. Next, we automatically identify relevant videos by extracting key-frames (frames containing significant visual changes from previous frames) from a video and using pre-trained classifiers to classify each key-frame. We set specific thresholds per-domain for minimum amount of visual change required to trigger a key-frame, since each modality is intrinsically different e.g. X-rays tend to be static images, while ultrasounds tend to be short repeating video clips (See section 7.2 in the Appendix). Finally, we identify narrative-style videos. For static domains (CT, X-ray, MRI), we select key-frames predicted to be medical and extract the next three (3) key-frames, computing the cosine similarity between the selected frame and each of the subsequent three frames. If all three have similarity scores $\geq$ a preset threshold of $0.9$, we count it as a narrative streak. A video is tagged as narrative if a domain-specific preset percentage (%) of the selected frames exhibit a narrative streak. For non-static domains like ultrasound, we utilize the video’s transcript to look for speech around key-frames. A video is considered narrative if more than half the key-frames have text for more than a domain-specific number of seconds. Consequently, we download all narrative-style videos at full resolution.
Articles: Since these articles/studies are carefully written and reviewed, we inherit the benefit of this data source being inherently filtered and well curated by authors and reviewers of the constituent journals and simply classify each figure into one of the 12 medical imaging domains.
3.2.3 Text Extraction and Denoising.
Videos: In line with Quilt-1M [54] we leverage an open-source ASR model - Whisper [103] to transcribe all speech from the selected videos, making sure to account for transcription errors using a similar methodology of finding these types of errors and correcting with a language model. See section 7.4 in the appendix for further details of the error-extracting algorithm, medical n-gram verifier, and methodology behind leveraging a general-purpose LM for error correction.
Articles: Similarly we parse each article’s XML document, extracting each figure’s caption and inline mentions. Since many sub-figures are typically grouped into single large figures, we split the compounded figure captions into sub-captions, leveraging a language model to find and split sub-captions due to the non-triviality of identifying enumerations in the text and splitting the captions. Furthermore, we refine the inline mentions of a figure and match them to specific sub-captions/sub-figures; please see more details of this pairing in sections 8.1 8.4 and 8.6.
3.2.4 Image Extraction and Denoising.
Videos: For each video, we identify medical key-frames and subsequently leverage these frames’ times to split the video into time-intervals called chunks from which to extract representative image(s). To extract representative image(s), we use the median image (pixel-space) of stable (static) frames in each chunk if they exist, else we de-duplicate the captured key-frames, exploiting the human tendency in pedagogy videos to pause while explaining and pointing [100, 46, 118].
Articles: For scientific documents, we extract the figures as images. However, many of these figures contain multiple sub-figures, which can take nonconventional grid shapes and are labeled irregularly making the task of splitting into sub-figures and pairing with the correct sub-caption non-trivial. Since most compound figure layouts are not uniform and vary in the whitespace in between sub-figures, we train an object detection model based on the YOLO architecture [59] on sub-figure annotation datasets MedICaT and ImageCLEF 2016 [122, 38]. See more details in section 8.3 of the appendix including examples of non-trivial cases.
Figure 3: Zeroshot Classification Results shows that our model GenMedCLIP outperforms all other baselines including the out-of-domain CLIP, and biomedical vision-language models BiomedCLIP, and PubMedCLIP across the constructed medical benchmark which covers all 11 medical domains represented. The metric for Xray and Mammography is mean average precision while the rest is accuracy.
3.2.5 Extracting Localizing Traces
Extracting the narrative traces/cursor location from medical clips poses a significant challenge due certain in-domain issues including homogeneity in color and texture, significant black/white background, alongside the presence of other dynamic elements in videos, such as minor pixel variations and variations in the narrators’ cursor movement speed and style; hence, we adopt the methodology proposed by Quilt-LLaVA [118] centered around the observation that narrators typically pause before signaling with their cursor, guiding the audience’s attention to specific relevant regions while describing those regions. This allows us to isolate segments in each video where the background is mostly static, termed stable chunks. To detect the stable chunks, we develop a frame-differencing approach to detect chunks with minimal background movement. Our algorithm computes the absolute difference between consecutive frames and then applies a Gaussian filter for adaptive thresholding to detect frames with minor changes.
Given the homogeneity in medical images and backgrounds, this naive pixel-wise difference algorithm yields a high rate of false positives, namely, chunks that had a change occur in them are identified as stable. To alleviate this, we incorporate a perceptual metric, by evaluating the structural similarity index measure (SSIM) on randomly sampled patches to further verify whether a change has occurred between two frames.
Next, for each stable chunk where the background remains consistent, we calculate a median frame in the pixel domain to represent the background. This median frame is then subtracted from the frames within the stable chunk, and a threshold is applied to mitigate noise or other artifacts, effectively isolating the cursor. Subsequently, we capture the cursor by identifying the coordinates of the maximum pixel value. However, this approach assumes the absence of other slight movements with respect to pixel changes, which is not always the case, as there may be movements like the narrator’s facial expressions within the scene. We address this by applying a face detection model [115] to mask such distractions, ensuring our analysis concentrates solely on cursor movement. This algorithm offers a surprisingly generalizable way to collect cursor traces from any medical video in static domains like X-rays and histopathology. Finally, we extract the smallest bounding boxes around the traces.
3.3 Aligning modalities
Videos: Following prior work [54], to align image, text, and trace modalities we compute time chunks for each video denoted as $[(t_{1},t_{2}),(t_{3},t_{4}),\cdots(t_{n-1},t_{n})]$ from key-frames after discriminating for medical frames using the pretrained classifiers – (scene_chunks). Each scene_chunk is padded with pad_time to its left and right. We use the methods described above to extract the medical/ROI captions as well as the representative image(s) for every chunk/time-interval in scene_chunks Finally, each chunk in scene_chunks is mapped to text (both medical and ROI captions), traces and images. Next, we map each medical image to one or more medical text (with traces). Using the time interval in which the image occurs, we extract its raw text from ASR, then extract and correct keywords using the Rake method, which we refer to as raw_keywords. We extract keywords from each medical text returned using the LLM, and we refer to these as keywords. Finally, if the raw_keywords occur before or slightly after a selected representative image, and overlap with the keywords in one of the Medical/ROI texts for that chunk, we map the image to the medical/ROI text. Traces are encoded as the cartesian position of the cursor relative to the image size, we use $(x^{t}_{j},y^{t}_{j})$, where $x\in[0,W]$ and $y\in[0,H]$, with $W$ and $H$ representing the image width and height, respectively, $t$ spans from 0 up to the total duration of the $j^{th}$ stable chunk.
Articles: The majority of our curated PubMed data uses alphabetic labels in compound figures to denote sub-figures, which increases the complexity of pairing individual sub-figures from compound figures to sub-captions. Our solution leverages an optical character recognition (OCR)  [^1] https://github.com/JaidedAI/EasyOCR on each sub-figure to detect the sub-figure labels, which we then match to the extracted sub-caption labels. To improve the reliability of this approach, we upscale the detected sub-figures by a factor of 4 to enlarge the sub-figure text label for OCR. We also impose a 95% confidence threshold on predicted text to isolate the sub-figure label, as text detected at lower confidence is often non-label text present in the figure (e.g. axis titles, graphs). We then attempt to match the detected sub-figure label with the sub-caption label. If a match is found, we pair the selected sub-figure and sub-caption. Despite the generality of this approach, we identified a few failure cases and proposed an error handling solution for these cases in section 8.5 of the appendix.
3.4 Characterizing MedicalNarratives
To create MedicalNarratives we combine medical narratives curated from videos with image-text pairs curated from PubMed, resulting in 4.7M total image-text samples of which 1M samples are localized narratives.
3.4.1 Narratives
We searched over 738K videos and extracted 74K narrative-styled videos that passed our heuristics and had relevant medical imaging pedagogy, a $10.1$% yield making up a total of 4526 hours of video. In total, we collect 809K unique images with an average size of H: 1521px and W: 903px and 1.62M image-text pairs from videos with 1M of these samples grounded with traces, these samples cover 101.6M number of unique trace points yielding 546K number of unique bounding boxes with an average size of H: 291px and W: 357px across the 4 domains with traces: CT, MRI, X-ray, and Histopathology. The mean length of the text captions is 22.37 words, with an average of 1.88 medical sentences per image. Our dataset spans over 4M UMLS entities from those mentioned in the text with over 300K unique entities across medical (e.g., findings, or disease) and non-medical (e.g., governmental or regulatory activity) semantic types.
3.4.2 Non-Narratives
We extract 5.4M articles from PubMed [90], with 23M figures, after filtering for medical figures only, we obtain 1.03M figures from 273K articles, and after sub-figure separation, we have an average of 2.62 subfigure-subcaption pairs per-article figure, with an average of 45.45 words per-caption.
3.5 Quality
Unlike localized narratives [100, 128] where localization accuracy can be measured by comparing against human annotation, none of our videos to our knowledge have any structured human spatial annotation to compare against. Nonetheless, to evaluate our pipeline’s performance, we assess several aspects. First, we calculate the precision of our LLM’s corrections by dividing the number of conditioned misspelled errors replaced (i.e., passed the UMLS check) by the total number of conditioned misspelled words found, yielding an average of 47.99%. We also determined the unconditioned precision of the LLM, similar to the previous step, and found it to be 17.58%. Therefore, we replace our detected incorrect words with the LLM’s correction 47.99% of the time, and 17.58% of the time we replace the LLM’s detected errors with its correction. To estimate the ASR model’s transcription performance, we compute the total number of errors replaced (both conditioned and unconditioned) and divide it by the total number of words in each video, resulting in an average ASR error rate of 0.81%. Also note that, by prompting the LLM to extract only medically relevant text, we further eliminate identifiable information, such as clinic addresses, from our dataset.
4 GenMedCLIP: Experiments and Results
Since MedicalNarratives supports semantic tasks, we test the utility of our dataset on two medically relevant tasks image classification (zeroshot and linear probing) and cross-modal information retrieval (zero-shot) across all in-domain modalities. We select the Contrastive Language-Image Pre-training (CLIP) objective [102] to pre-train a VL model we call GenMedCLIP. We train several models, varying the image and text encoders and making adaptations inline with prior work on the choice of encoders and text tokenization for improved performance [140, 54]. For the image tower we finetune Vision Transformers (ViT-Base) [34] models pretrained using a supervised cross-entropy objective (ViT-Base-16 and ViT-Base-32 [132]) and unsupervised contrastive objective ViT-Base-16) [102], on 224*224 pixel images. On the text tower, we use GPT2 [101] with a context length of 77, and BioMedBert [45] with context size to 256. To train our models we utilize the open-source implementation of CLIP [102] called OpenClip [55] on 4 Nvidia A40 GPUs for 20 iterations over the entire train dataset of  4.2M samples. To ensure a fair comparison with baselines we trained three different variants of our model: GenMedCLIP-32: with ViT-B/32 image-tower with GPT2/77 text-tower architecture, GenMedCLIP-PMB: with ViT-B/16 image-tower and Bert/256 BiomedBert [45] text-tower architecture, and GenMedCLIP-PMB: with ViT-B/16 image-tower with GPT2/77 [45] text-tower architecture; all finetuned for 20 epochs over our train-set. Please see training hyperparameters used in Section 10 of the Appendix.
Model Isic Til Pcam Mhist Nck Mammo Avg
CLIP-ViT-B-16 [102] 71.23 91.23 82.42 63.97 92.26 83.30 80.74
PubMedCLIP [36] 68.58 91.32 84.07 72.16 92.29 83.90 82.06
BiomedCLIP [140] 68.25 91.82 83.43 66.73 93.05 83.70 81.17
GenMedCLIP-32 72.75 93.26 86.77 72.06 92.77 83.70 83.55
GenMedCLIP-PMB 69.38 91.51 84.54 67.66 88.02 84.20 80.88
GenMedCLIP 74.87 93.34 87.69 72.16 90.84 84.90 83.97
Table 1: Linear Probing results across datasets representing Dermatology (Isic), Histopathology (pcam, mhist, nck), and Mammography (vinDr-Mammo) classification tasks. GenMedCLIP outperforms all baselines showing the capacity of our model to be fine-tuned for downstream tasks. The metric used is accuracy.
4.1 Benchmarking on Downstream Medical Tasks
We evaluate the utility of GenMedCLIP on a new medical imaging benchmark of all medical domains represented in our pre-training dataset MedicalNarratives, with some domains represented by $>=1$ dataset/task for classification, totaling 29 downstream datasets and on a held-out set of 1000 unique images for the retrieval task downstream. For MRI we use the RadImageNet [83] MRI subsets tasks based on the anatomical region scanned in the image these include Ankle/foot with 25 classes, Brain with 10 classes, Knee with 18 classes, Abdomen/pelvis with 26 classes, Hip with 14 classes, Shoulder with 14 classes, Spine with 9 classes. To evaluate on CT domain we also use RadImageNet’s [83] CT dataset which cover two (2) anatomical regions with Lung having 6 sub-classes and Abdomen/pelvis with 28 subclasses. For ultrasound, we evaluate on RadImageNet’s [83] US dataset which covers a total of 15 classes across Thyroid and Abdomen/pelvis anatomical regions. For Xray, we evaluate on VinDr-CXR Chest Xrays [91] test set and report the mean average precision (mAP) across all 28 findings, similarly to evaluate on Mammography we use VinDr-Mammo [92] and report the mAP on all X findings, leveraging only the standard bilateral craniocaudal (CC) view of the test set. We evaluate on surgical organ classification using Dresden [22] which covers 8 abdominal organs; to evaluate for endoscopy domain we test on all procedures images in GastroVison [56] with 27 classes. For Dermatology we evaluate on the Diverse Dermatology Images (DDI) [31] binary (benign or malignant) dataset and Isic 2018 dataset [28]. For Dentistry we evaluate on Dental orthopantomography (OPG) [104] X-ray dataset with 6 classes. To evaluate the Ophthalmology domain we evaluate on G1020 [13] a retinal fundus glaucoma dataset and on Optical Coherence Tomography Dataset (OCTDL) [72] with 6 disease classes. We evaluate the Histopathology domain on the following datasets: PatchCamelyon [127] for lymph node metastatic tissue binary prediction task, NCT-CRC-HE-100K [63] on 8 morphological classes, BACH [10] which consists of breast tissues with 4 classes including being and invasive carcinoma, Osteo [12] osteosarcoma dataset with 3 classes including necrotic tumor, SkinCancer [70] dataset of tissue patches from skin biopsies of 12 anatomical classes and 4 neoplasm categories that make up the SkinTumor Subset, we also evaluate on MHIST [131] dataset of colorectal polyps tissue, LC25000 [18] dataset, which is split in-between LC25000 (Lung) and LC25000 (Colon), for lung and colon adenocarcinomas classification, and on TCGA-TIL [113] for tumor-infiltrating lymphocytes (TILs) binary classification, based on H&E images from 13 of The Cancer Genome Atlas (TCGA) tumor types.
Models T2I retrieval I2T retrieval Avg
@5 @50 @200 @5 @50 @200
CLIP-ViT-B-16 [102] 3.48 20.38 35.69 3.56 20.39 35.42 19.82
PubMedCLIP [36] 1.44 12.68 25.44 1.10 12.30 24.07 12.84
BiomedCLIP [140] 16.50 51.48 67.46 15.71 48.85 64.61 44.10
GenMedCLIP-32 22.36 76.33 88.60 20.75 75.15 88.23 61.90
GenMedCLIP-PMB 28.29 82.91 92.43 29.21 82.91 92.43 68.03
GenMedCLIP 34.89 83.83 92.27 34.26 83.48 92.32 70.17
Table 2: Retrieval results on our held-out set of 16K samples across all medical domains, show that our model GenMedCLIP outperforms all other baselines on both Zeroshot image-to-text and vice-versa text-to-image retrieval task.
4.2 Zero-shot classification
A model must maintain a comprehensive understanding without requiring specific data for retraining. Thus, we evaluate our model’s zero-shot performance against three state-of-the-art models: CLIP, BiomedCLIP, and PubMedCLIP. In Figure 3, each domain in the benchmark is represented by a set of datasets (as provided above). The prompts used for these evaluations are presented in Table 6 in the Appendix. Across the benchmark our model average the following GenMedCLIP-32: 31.33%, GenMedCLIP-PMB: 31.46%, and GenMedCLIP: 32.55% metric all outperforming BiomedCLIP with 27.80 overall by a max value of 4.75%. Specifically, as shown in Figure Fig. 3, GenMedCLIP outperforms all baselines in 5 medical domains: Histopathology, Dermatology, Mammography, Surgery, Ultrasound, and CT. GenMedCLIP is also comparable to baselines in the Chest X-ray, Endoscopy, Ophthalmology, and MRI domains.
4.3 Supervised linear probing
We assess the full-shot performance of our model by conducting linear probing with 100% of the training data; we report the average accuracy over all benchmark evaluation across five distinct datasets, specifically those with dedicated training and testing sets among our external datasets in Dermatology, Histopathology, and Mammography. Remarkably, our model, utilizing the ViT-B/32 architecture with GPT/77, outperforms its counterparts, BiomedCLIP, and CLIP, in most datasets. Overall, on average GenMedCLIP outperforms all other models including BiomedCLIP and PubMedCLIP with over 2.8%, and over 1.9% respectively.
4.4 Cross-Modal Retrieval
We evaluate cross-modal retrieval performance by examining both zero-shot text-to-image and image-to-text retrieval capabilities. To do so we leverage a randomly selected held-out partition of MedicalNarratives, not used in training our models. The held-out set contains 16K image-text pairs with the following medical modality distribution: 1756 X-ray, 1237 MRI, 1851 CT, 1351 Ultrasound, 1744 Surgery, 1346 Endoscopy, 1189 Dermatology, 1216 Dentistry, 1151 Ophthalmology, 1000 Histopathology, 1299 General Medical, 1149 Other (Mammo etc) image-text pairs. We conduct retrieval, in our study, by identifying the nearest neighbors for each modality and then determining whether the corresponding pair is within the top N nearest neighbors, where N $\in$ 1, 50, 200, mimicking several medical search tasks. Results in Table 2 shows that on average GenMedCLIP outperforms all baselines and specifically outperforms BiomedCLIP by 26.07%, The results also confirm the observation made in BiomedCLIP [140] where the general CLIP model outperforms the in-domain model PubMedCLIP by 6.98%
5 Conclusion
In this work, we propose a robust protocol for annotating medical narratives, aiming to standardize the process of constructing multimodal medical datasets. Our curated dataset MedicalNarratives, which follows the Narratives Annotation Protocol addresses the specific challenges of medical data collection at scale balancing the relationship between utility and ease/cost of collection. We argue that this protocol can serve as the de facto standard for annotating future multimodal medical datasets, particularly given its flexibility in capturing grounded text describing medical images effectively. By establishing and scaling this protocol, we pave the way for systematic and scalable dataset creation for all medical domains.
For semantic tasks, we demonstrate a strong performance over prior vision-language (VL) models on our well-curated medical imaging benchmark, across both classification (zero-shot, and full-shot) and retrieval tasks, establishing new VL state-of-the-art results and demonstrating the effectiveness of in-domain data filtration methods on model performance, as we finetune our GenMedCLIP on 4.7M samples while BiomedCLIP trains on over 15M samples. We hope future work leverages our developed models, dataset, and protocol.
6 Limitations
While our approach and findings contribute significantly to the development of multimodal medical datasets and model performance in the medical domain, certain limitations need to be acknowledged:
  1. 1.
    Underlying Bias in Abnormality Representation: A major drawback in our dataset is its overrepresentation of abnormal cases, a bias that mirrors reality as in hospitals, diagnostic imaging is typically performed only after significant clinical suspicion of pathology. Consequently, our data disproportionately captures abnormal findings, which may affect the model’s generalizability to balanced populations and lead to bias in clinical decision support.
  2. 2.
    Focus on 2D Medical Imaging: Our work is limited to 2D medical images, despite the prevalence of 3D imaging modalities (such as MRI, CT, and OCT scans). While recent studies suggest that 2D approaches can sometimes rival or surpass native 3D approaches, there remain applications where 3D data could offer distinct advantages. Expanding our dataset and methodology to include 3D imaging would be an important step toward broader applicability and model performance.
  3. 3.
    Exclusion of Other Medical Data Modalities: This work currently focuses solely on medical imaging data, excluding other critical types of medical data such as laboratory tests, time-series signals, genetic data, and patient histories. These additional data sources are valuable for building comprehensive and holistic patient models, and their integration is key for achieving the goal of generalist medical models [144, 124].
  4. 4.
    Lack of Localization Accuracy: Our dataset does not contain direct human-annotated ground truth for localization, such as bounding boxes that indicate the specific regions associated with the text. This limitation impacts our ability to measure the overlap between trace annotations collected from videos and human annotations.
Supplementary Material
7 MedicalNarratives: YouTube Video Curation
Figure 4: The data curation pipeline for the PubMed subset of the MedicalNarratives dataset. Download: downloading PMC-OA. Extraction: extracting figures, captions, and inline references. Filtering: filtering for medical images. Sub-captions: splitting compound figure captions into sub-captions. Sub-figures: detecting and cropping sub-figures from compound figures. Pairing: matching sub-figures and sub-captions. Inline pairing: matching inline mentions of figures with the most relevant sub-figure or sub-figures.
Distilling the volume of data YouTube offers into a grounded vision-language dataset that captures the all available signal of medical pedagogy video data is a significant task. Each step in the data curation process presents unique challenges when scaling to handle multiple medical domains.
With MedicalNarratives, we collect vision-language datasets grounded in time with language-correlated traces across twelve medical domains with the first three domains defined to be static where representative samples are usually static images: (1) computed tomography (CT), (2) magnetic resonance imaging (MRI), and (3) xray, and non-static domain with representative samples exhibiting significant visual change: (4) ultrasound, (5) surgery, (6) endoscopy, (7) dentistry, (8) dermatology, (9) mammography, (10) ophthalmology, (11) histopathology and (12) general medical illustrations. When processing these subsets, our approach differs to accommodate the nuances of the video data. Our data curation pipeline can be split into these high-level tasks of:
  1. (A)
    Searching for representative videos in each medical domain.
  2. (B)
    Filtering videos for narrative style.
  3. (C)
    Extracting image, text, and cursor traces from selected videos.
  4. (D)
    Denoising and de-duplicating the collected raw data.
  5. (E)
    Aligning image, text, and localization traces.
  6. (F)
    Collecting metadata useful for varying downstream tasks (e.g. subdomains) and interleaving the dataset.
In the following sections, we present a detailed overview of the major steps in curating MedicalNarratives starting with search. We also present examples of curated narrative samples in Figures 5, 6, 7, and 8.
Figure 5: X-ray examples within the MedicalNarratives dataset. Input: raw input text from ASR. Output: the output from the LLM, with denoised medical and ROI text. Traces: Cursor traces and bounding boxes aligned in-time with the raw text. UMLS: UMLS entities extracted from the medical text. Subdomain: classification of the sample into finer-grained subdomains.
Figure 6: Histopathology examples within the MedicalNarratives dataset. Input: raw input text from ASR. Output: the output from the LLM, with denoised medical and ROI text. Traces: Cursor traces and bounding boxes aligned in-time with the raw text. UMLS: UMLS entities extracted from the medical text. Subdomain: classification of the sample into finer-grained subdomains.
7.1 Domain-Specific Search
We first identify medical channels and videos for each domain on YouTube, using keywords from online medical glossaries specific to each imaging modality or medical domain. To increase the percentage of narrative or educational style videos, a list of priority keywords: "educational", "interpretation", "case study", and similar phrases, are appended to search keywords. We limit our search to channels with <1M subscribers since some channels focus on multiple domains (e.g. radiology channels span CT, MRI, X-ray) and therefore might have a large subscriber base, and channels with >1M subscribers often contain non-imaging videos.
We observe during channel search that searching YouTube for channels by keyword tends to produce irrelevant results, hence, we adopt a video-first search strategy: since video titles are more informative than channel titles, we first find relevant videos, then evaluate the channel of the relevant video for more hits. Each video result is downloaded in low resolution for further analysis. To limit searching irrelevant channels we implement early stopping, wherein, if the first 10 videos of a channel fail the medical filtering step, the channel is skipped, allowing us to keep compute cost low while increasing our pool of visited channels.
7.2 Medical Filtering
Each potential pedagogy video is evaluated by the following heuristics:
  1. 1.
    The video duration is longer than 1 minute and shorter than 2 hours as videos outside this range usually contain little medical imaging information.
  2. 2.
    Video contains speech. We check this either through the video’s transcript from the YouTube API, or if not present using the inaSpeechSegmenter  [^2] https://github.com/ina-foss/inaSpeechSegmenter tool on the first minute of audio.
  3. 3.
    The number of medical scene frames exceeds the empirically determined threshold unique to each medical domain. This heuristic filters for narrative-style videos (See Section 7.3).
To expand on the third heuristic, we extract the key-frames of a video for classification; for static domains, we utilize FFmpeg  [^3] https://github.com/FFmpeg/FFmpeg to detect scenes and extract key-frames (frames with significant visual changes from previous frames). We experiment with scene detection thresholds to determine the optimal threshold per domain across various video durations. For non-static domains, we leverage adaptive content scene detection in  [^4] https://github.com/Breakthrough/PySceneDetect to avoid capturing frames that are visually different but still part of the same shot (which are characteristic of non-static domains). Camera movements are common in domains such as surgery or endoscopy, and nearly duplicate frames that would be generated by thresholding on video content are instead merged when using PySceneDetect’s adaptive detection algorithm. We specifically tune the adaptive detection for each domain by experimentally determining parameters for the algorithm on sample videos from each domain.
We then classify the key-frames of a video using pre-trained classifiers per domain (see Section 10.3). Using the percentage of key-frames predicted to be medical images, videos are differentiated into three categories: positive videos, near-positive videos, and negative videos. For example, a video with 50% of key-frames predicted to be MRI images is a candidate for further processing, while a video with only 2% of key-frames is not. Positive videos contain sufficient medical content for the given domain, while near-positive videos may or may not contain sufficient medical content. The thresholds defining positive/near-positive/negative are unique to each domain. We then manually examine a subset of the near-positive category, and determine a more fine-grained percentage threshold to extract more positive videos out of the pool of near-positive videos. See Table 3 for the final percentage thresholds used.
Domain Threshold (%)
CT & X-ray 10
MRI 5
Dermatology & Dentistry 30
Endoscopy & Surgery 50
Ultrasound 40
Ophthalmology 35
Mammography 25
General medical illus. 20
Table 3: Final percentage thresholds used during video key-frame classification.
7.3 Narrative Filtering
We define narrative-style videos as pedagogy videos where the narrator focuses on describing or analyzing medical images onscreen. To filter for these videos, we first check the first minute of each medical video for speech using inaSpeechSegmenter to ascertain the presence of a narrator.
For static domains like X-ray, we define a narrative streak as any partition of the video where frames sampled close (w.r.t. time) together are similar using cosine similarity, indicating the narrator is spending a lot of time analyzing that frame. Specifically, we randomly sample a fixed number of clips across each video, sampling three consecutive frames from each clip and checking for similarity. Finally, if the number of streaks in each video is greater or equal to the set threshold for that domain, the video is considered to be narrative. This simple filtering algorithm helps us reduce the number of videos we process from 748k to 74k videos.
For non-static domains like surgery or ultrasound, consecutive key-frames often exhibit significant change so we instead look for persistent narration around key-frames classified as medical. For example, for ultrasound clips, we extract the times for each consecutive positive key-frame accumulating a sequence of start and end times. Within these time intervals, we determine whether speech exists either through the video’s YouTube API transcript or by extracting the audio during the selected time interval and using inaSpeechSegmenter to determine if the segment contains any speech.
7.4 Text Extraction using ASR and Text Denoising.
In line with Quilt-1M [54] we leverage an open-source ASR model - Whisper [103] to transcribe all speech from the selected videos and make sure to account for transcription errors using a similar methodology of finding these types of errors and correcting with a language model. We use the whisper-large-v2 model in the stable-ts library for word-level and sentence-level transcription. As anticipated, this model often misinterprets medical terms, thus requiring the use of post-processing algorithms to minimize its error rates. For this, we adopt a similar methodology proposed in Quilt-1M [54] to identify, correct, and verify these errors, please see section A.1 in Quilt-1M [54] supplementary for more details.
From the ASR-corrected text, we extract medical text which describes the image(s) as a whole. Also, when the speaker describes/gestures at visual regions-of-interest through statements like "look here …", we extract the text entity being described as ROI text in line with Quilt-1M [54].
To extract relevant text, we prompt LLMs to filter out all non-medically relevant text, providing context as necessary, while conditioning the LLMs to refrain from introducing new words beyond the corrected noisy text and set the model’s temperature to zero. Lastly, the LLM is used to categorize our videos into subdomains by conditioning with a few examples and prompting with the corrected video transcript as input (see Figure 11 for prompt and sample input/output).
Hyperparameter Training
Batch size (per GPU) 256
Epochs 20
Peak learning rate 1e-5
Learning rate schedule cosine decay
Warmup (in steps) 2000
Augmentation Resize; RandomCrop (0.8, 1.0)
Optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.98
Weight decay 0.2
Optimizer AdamW
Table 4: Training hyperparameters for GenMedCLIP
Hyperparameter ResNet50 ViT-Small
Batch size (per GPU) 256 32
Epochs 10 100
Peak learning rate 1e-2 1e-3
Learning rate schedule - cosine annealing
Augmentation RandomResizedCrop RandomResizedCrop
(224), (384, 0.98, 1.0),
Resize RandomHorizontalFlip
Optimizer momentum 0.9 0.9
Weight decay 1e-4 -
Optimizer SGD SGD
Table 5: Training hyperparameters for domain classifiers.
Datasets prompt
Pcam [127], Nck [63], Lc2500 [18], Mhist [131] [ "a histopathology slide showing {c}.", "histopathology image of {c}.", "pathology tissue showing {c}.", "presence of {c} tissue on image" ]
Bach [10], Skin [70], Osteo [12]
Tcga_til [113], DDI [31], Isic [28], Dental [104] [ "{c} presented in image", "evidence of {c} in image", "an image showing {c}" ]
Gastrovision [56], G1020 [13], Octdl [72], VinDrM [92]
VinDrXR [91], Dresden [22] , Radimagenet [83]
Table 6: Zero-shot classification templates used to evaluate GenMedCLIP’s zero-shot capacity across all multiple dataset that constitute the medical benchmark.
Models Xray (mAP) CT MRI Mammo (mAP) US Optha Derm Endo Surg Dental Histo Overall
vindrXR lung abd af brain hip knee abd shdr spine vindrM rad g1020 octdl ddi isic gastro dresden dental til pcam lc_lung nck skin skin_tumor lc_colon mhist bach osteo
CLIP-ViT-B-32 6.95 2.74 4.14 2.75 0.77 2.67 1.86 2.23 3.13 12.29 10.64 8.45 69.61 10.22 41.31 21.76 4.94 10.66 - 21.33 61.81 61.55 29.20 4.47 9.84 65.57 50.67 25.25 58.85 21.63
CLIP-ViT-B-16 6.55 1.61 3.76 3.59 6.55 3.00 1.13 1.49 4.29 14.07 10.65 5.41 31.47 6.98 60.67 13.23 1.77 11.38 2.51 20.32 51.80 47.06 21.41 5.55 13.22 79.56 52.61 23.75 43.54 18.93
PubMedCLIP 7.40 6.06 2.04 1.08 8.95 1.38 4.07 0.54 8.50 31.21 10.45 15.94 68.33 7.56 35.52 18.19 2.05 9.72 12.19 24.55 50.38 33.33 26.45 8.07 23.01 63.66 62.74 15.25 43.63 20.77
BiomedCLIP 10.29 6.80 2.53 3.11 12.31 2.98 4.88 0.97 6.09 11.35 10.57 58.10 29.12 18.60 51.22 20.70 3.49 16.37 16.83 37.03 71.71 71.34 49.17 24.83 40.39 84.98 38.59 44.25 46.75 27.43
GenMedCLIP-32 10.23 14.25 1.59 2.25 23.40 3.68 9.45 1.18 2.36 24.00 10.30 44.98 66.67 20.35 62.96 19.78 2.10 15.04 26.89 23.14 70.56 81.11 48.05 28.45 49.67 93.24 55.37 37.25 55.82 31.18
GenMedCLIP-PMB 9.90 10.36 2.30 3.95 8.80 1.80 4.33 0.98 8.38 25.04 12.06 52.98 29.22 24.56 57.01 37.63 2.08 19.30 16.83 49.63 71.90 82.05 51.05 39.32 48.34 71.68 61.82 52.00 42.44 30.96
GenMedCLIP 9.66 27.35 1.38 2.75 7.52 2.61 9.93 2.80 3.10 22.59 11.12 63.48 33.53 21.22 72.26 37.10 2.38 18.97 16.44 20.34 65.90 72.37 52.16 42.37 59.87 94.16 60.59 52.50 41.43 31.99
Table 7: Expanded Zeroshot Classification Results shows that our model GenMedCLIP outperforms all other baselines including the out-of-domain CLIP and biomedical vision-language models BiomedCLIP and PubMedCLIP across the constructed medical benchmark. The benchmark covers all 11 medical domains represented, excluding the non-medical domain of medical illustrations. The metric for X-ray and Mammography is mean average precision while the rest is accuracy.
Domain CT MRI X-ray
Image-text-ROI pairs 34224 41019 32241
Image-text-ROI-text pairs 51041 54213 49639
Avg. ROI Text/Image ROI 3.29 2.9 3.82
Num. ROI Text/Video 96771.0 84409.0 85682.0
Avg. Words/ROI Text 10.65 9.61 12.32
Avg. ROI UMLS/Text 1.47 1.48 1.47
Avg. ROI/Image 1.61 1.38 1.74
Avg. ROI Text/Chunk 2.61 2.33 2.54
Unique ROI BBox 45465 34709 49208
Unique ROI Traces 11386480 4744770 10650027
Avg. ROI Chunk Duration 12.94 6.23 14.39
Avg. BBox Height 321.7 205.1 312.44
Avg. Bbox Width 543.12 282.65 507.58
Table 8: Characterization of MedicalNarratives image-text-trace subset, categorized by individual medical domains. The table provides detailed statistics for each medical modality, including the number of unique images, total dataset duration, ASR error rate, and average image resolution. Note: "ROI" in the table is shorthand for traces.
7.5 Identifying Cross-Modality Samples
To aid in interleaving the dataset we find videos with multiple medical modalities by leveraging a language model conditioned with a prompt with a list of all relevant medical modalities, the transcript of the video, 3 human-curated examples, prompting the model to output the list of modalities that are relevant to the transcript given (see Figure 10). Finally, we run the respective domain classifiers on each of the representative images in the video to determine which images (samples) are relevant to each identified domain.
8 MedicalNarratives: PubMed Article Curation
To curate the PubMed subset of MedicalNarratives, we download the PubMed Central Open Access Subset (PMC-OA) [90], containing 5.47 million articles and filter article figures for the same twelve domains as the YouTube subset of MedicalNarratives. Our data curation pipeline for PubMed is as follows:
  1. (A)
    Downloading PMC-OA and extracting each article’s XML and images.
  2. (B)
    Parsing each XML to extract figure captions and inline mentions of figures.
  3. (C)
    Filtering for figures with medical imaging with pretrained classifiers.
  4. (D)
    Splitting compounded figures and captions using fine-tuned object detection models and a language model.
  5. (E)
    Pairing correctly split sub-captions and sub-figures together using a combination of optical character recognition (OCR), bounding box heuristics, and error correction.
  6. (F)
    Matching inline mentions of figures with sub-figure/sub-caption pairs using a language model.
In the following sections, we will discuss the MedicalNarratives PubMed data pipeline.
8.1 Caption Extraction
From each obtained PubMed article, we extract the XML and image files for figure processing. The figure captions are extracted from the paper XML, cleaned, and paired with the corresponding image file of the figure. Additionally, we find all inline mentions of the figure and save them to the figure-caption sample. This yields 23.6M figure-caption samples.
8.2 Medical Filtering
To determine whether a figure belongs to one of the twelve domains of MedicalNarratives, we train a ResNet-50 CNN for binary classification. We use the same training datasets (see Table 10) selected when curating MedicalNarratives YouTube data, with a binary medical/non-medical label as the target prediction. This filtering step reduces the number of potential figures to 1.03M figures. To determine the specific domain or domains of each figure, we re-use the medical domain classifiers from the medical filtering step.
8.3 Sub-figure Detection
Figure 7: MRI examples with in the MedicalNarratives dataset. Input: raw input text from ASR. Output: the output from the LLM, with denoised medical and ROI text. Traces: Cursor traces and bounding boxes aligned in-time with the raw text. UMLS: UMLS entities extracted from the medical text. Subdomain: classification of the sample into finer-grained subdomains.
Figure 8: CT examples with in the MedicalNarratives dataset. Input: raw input text from ASR. Output: the output from the LLM, with denoised medical and ROI text. Traces: Cursor traces and bounding boxes aligned in-time with the raw text. UMLS: UMLS entities extracted from the medical text. Subdomain: classification of the sample into finer-grained subdomains.
The majority of figures after medical filtering are compound figures, which compress detailed information into a single image and caption. Splitting these compound figures into sub-figures is a non-trivial task, since there is no uniform compound figure layout. In contrast to Quilt-1M’s [54] image processing-based approach to splitting these figures, we opt for an object detection approach, which we empirically determined is capable of handling wider range of abstract layouts.
Specifically, we finetune a YOLO object detection model [59] to detect sub-figures within compound figures using two medical subfigure separation datasets: MedICaT’s sub-figure annotations and ImageCLEF 2016’s Figure Separation medical task [122, 38]. MedICaT contains 7507 sub-figure bounding box annotations from 2069 compound figures. ImageCLEF 2016 Figure Separation contains 6782 sub-figure bounding box annotations. We fine-tune a YOLOv8-Large [59] for 100 epochs using an 80/20 training/validation split on the subfigure separation data. One major advantage of the object detection approach is that our model can successfully detect sub-figures even when there is little to no gap/whitespace in-between sub-figures. We process each figure with the fine-tuned YOLOv8 sub-figure detection model, where each detected sub-figure is cropped, and up-scaled by a factor of 4. In the case of compounded figures with uncompounded caption i.e. all constituting images communicate a singular concept (see Figure 9) we pair the caption to the original compounded figure.
Figure 9: (a), (b) Compound figures successfully separated into sub-figures (green), which are then up-scaled and saved. (c) A figure that is incorrectly identified as a compound figure. Since the figure caption contains no sub-captions, the original figure will paired with the entire caption during sub-figure/sub-caption pairing.
8.4 Sub-caption Separation
A compound figure caption usually contains multiple sub-captions. A heuristics-based approach to splitting these compounded captions is difficult to design since figure sub-captions are labeled differently with article authors adhering to varying writing styles typically set by the publishing journal. We therefore opt for an LLM-based approach where we provide diverse examples of sub-caption separation, instructing the language model (GPT-3.5 Turbo) to follow the process below:
  1. 1.
    Separate the figure caption into sub-captions based on the sub-figure labels present in the caption e.g. "(A)", "I)", "a.", "bottom left", etc.
  2. 2.
    Strip the sub-figure labels from each sub-caption text produced.
  3. 3.
    If any context in the caption pertains to the entire figure, add this context to each sub-caption. This step ensures that each individual sub-caption retains the entire context of the figure.
  4. 4.
    Return each sub-caption paired with its sub-figure label.
We condition the LLM with a few examples, including handling non-compound figures and captions that use spatial cues (e.g. left, center, right) to refer to sub-figures (see full prompt and sample input/output in Figure 12). We also process the sub-figure labels returned from the LLM, stripping parentheses and other extraneous characters to make sub-figure/sub-caption pairing easier.
8.5 Pairing Sub-figures to Sub-captions
Given the separated sub-captions and sub-figures for a compound figure, next we tackle the problem of pairing the correct sub-caption with the correct subfigure. The majority of our curated PubMed data uses alphabetic labels in compound figures to denote sub-figures. Our approach therefore leverages optical character recognition on each sub-figure to detect the sub-figure labels, which we then match to the sub-caption labels extracted during section 8.4.
During the sub-figure detection step, we upscale the detected sub-figures by a factor of 4 to enlarge the sub-figure text label for OCR. We impose a 95% confidence threshold on predicted text during OCR to isolate the sub-figure label. Text detected at lower confidence is often other text in the figure (e.g. axis titles, graphs) being present. We then attempt to match the detected sub-figure label with the sub-caption label. If a match is found, we pair the selected sub-figure and sub-caption.
There are several types of cases where this approach requires error handling, e.g.:
  1. 1.
    In a single sub-figure, no labels are identified that exceed the 95% confidence threshold.
  2. 2.
    Sub-captions use spatial cues to identify sub-figures, e.g. "upper left", "center", "right".
  3. 3.
    If the number of detected sub-figures does not match the number of sub-captions: either some sub-figures or some sub-captions are unpaired.
In case 1, if the compound figure has exactly one sub-figure and one sub-caption left unpaired, we pair the two. Otherwise, we lower the confidence threshold to 80% and re-detect sub-figure labels, then re-match with sub-captions. Sub-figures that fall in this category tend to have their label close to the border of the cropped sub-figure, have small sub-figure text, or have backgrounds that resemble the font color of the label. For case 2, we use the bounding box coordinates of the detected sub-figures and the spatial cues provided in the caption to pair figures and captions. For example, a sub-caption with the label "upper left" will be paired with the sub-figure with the upper leftmost bounding box. Lastly, case 3 occurs when either sub-figure detection and/or sub-caption separation perform incorrectly. The majority of figures in this category occur when sub-figure detection identifies multiple sub-figures, but the figure caption contains no sub-captions. In this case, we pair the original figure and caption.
Domain CT MRI Endo Genmed Surgery Optha Mammo Derma Ultrasound X-ray Dentist
Unique images 45889 54093 43230 23836 74435 758 288 19627 39084 54714 15369
Image-text pairs 37648 47211 40027 20318 64567 284 42 12057 31138 42521 11552
Total Med UMLS 290143 254279 504753 222699 649566 3243 251 106291 319199 270256 57086
Avg. Med Text/Image 2.44 2.14 3.49 2.79 2.85 2.48 1.4 2.02 2.69 2.56 2.54
Num. Med Text/Video 78931 70733 134385 54251 183259 681 59 26904 83768 71794 14867
Avg. Words/Med Text 28.38 24.03 40.3 36.03 32.1 36.74 15.62 23.49 31.64 30.07 30.61
Avg. Med UMLS/Text 3.75 3.68 3.83 4.28 3.57 6.07 4.25 3.89 3.93 3.83 3.87
Total Chunks 40883 38462 42770 23800 73314 758 283 19420 38948 37056 7655
Avg. Chunk Duration 31.33 18.36 47.08 36.4 28.83 74.6 2.42 23.6 31.33 34.05 50.38
Avg. Med Text/Chunk 2.13 1.97 3.23 2.35 2.48 1.27 0.16 1.0 2.08 2.07 1.7
Avg. Images/Chunk 1.25 1.52 1.01 1.0 1.04 1.0 1.02 1.01 1.0 1.75 2.04
Avg. Image-Text/Chunk 1.07 1.26 0.93 0.83 0.88 0.49 0.12 0.47 0.76 1.33 1.23
Precision (Unconditioned) 0.16 0.15 0.18 0.17 0.18 0.2 0.33 0.23 0.16 0.19 0.2
Precision (Conditioned) 0.48 0.45 0.48 0.56 0.54 0.44 0.73 0.43 0.41 0.46 0.42
Clinical ASR Error Rate 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
Total Duration (hrs) 324.0 410.0 428.0 278.0 383.0 15.0 0.0 187.0 735.0 562.0 140.0
Avg. Duration (mins) 13.31 12.04 6.24 9.72 7.58 6.94 0.11 8.06 12.43 19.22 13.69
Total ASR len. (words) 2342164 3314197 2721970 2061495 2933993 89011 4399 1203913 6245504 4485366 1149694
Avg. ASR len. (words) 1602.03 1619.05 660.99 1198.54 965.45 659.34 879.8 863.64 1759.3 2555.76 1872.47
Avg. Image Width 1530.82 1243.16 1302.91 1582.66 1707.22 988.71 1496.0 1717.6 1582.51 1491.5 1831.86
Avg. Image Height 954.37 789.68 759.47 1026.71 995.96 627.85 1128.0 998.99 910.32 923.71 1045.61
Table 9: Characterization of MedicalNarratives image-text subset, categorized by individual medical domains. The figure provides detailed statistics for each medical modality, including the number of unique images, total dataset duration, ASR error rate, and average image resolution.
8.6 Inline Figure Reference Pairing
In the final step of the pipeline, we pair the inline reference of a figure with the figure caption since inline references contain valuable context about the figure. However, an inline reference may refer to a sub-figure instead of the entire figure. We therefore utilize a language model to determine which sub-figure is most relevant to an inline reference. For each sample, we prompt an LLM (GPT-3.5 Turbo) with the list of sub-figure labels and a list of inline references and task the model with determining which sub-figure label best corresponds to the inline reference. In the case that the reference cites the entire figure instead of a sub-figure, we consider the inline reference relevant to all sub-figures. For each relevant sub-figure, we add the inline reference to its list of captions. See Figure 13 for the complete prompt and sample input/output.
9 Characterizing MedicalNarratives
To create MedicalNarratives we combine medical narratives curated from videos with image-text pairs curated from PubMed, resulting in 4.7M total image-text samples of which 1M samples are localized narratives. Section 3.4 gives an overview characterization of the entire dataset, and Tables 9, 8 below provide additional specific characterization details split per domain. Note we omit characterization for Histopathology in the tables below as the details for the domain can be found in prior work.
Figure 10: The GPT-3.5 Turbo prompts used to determine whether a video contains discussion of multiple medical domains, with few-shot examples.
Figure 11: The GPT-3.5 Turbo prompts used for determining which specific sub-domains are discussed in a video, with few-shot examples.
Figure 12: The GPT-3.5 Turbo prompts used for splitting a compound figure caption into sub-captions, with few-shot examples.
Figure 13: The GPT-3.5 Turbo prompts used for pairing inline references of a figure with the most relevant sub-figures, with few-shot examples.
10 Model Training and Evaluation
10.1 GenMedCLIP Training
We leverage OpenCLIP [55] to train our models as it allows us to quickly import our datasets and adapt varying components of our model including the underlying image and text towers and the training hyperparameters. Our experiments utilize Pytorch on 4 NVIDIA L40s GPUs, as well as gradient checkpointing, automatic mixed precision with bfloat16 to reduce memory usage. All other hyperparameters used are listed in Table 4. Our dataset is split into 16 tar files in the WebDataset  [^5] https://github.com/webdataset/webdataset format for training.
10.2 Evaluation
To evaluate zero-shot classification capacity across all constituting datasets in our medical benchmark outlined in 4 we leverage simple prompts listed in Table 6, with the specific results shown in Table 7.
10.3 Search Classifiers
To classify images into domains, we train a ResNet50 for 10 epochs and a ViT-Small for 100 epochs using DINO on a binary classification task for each medical domain. Both types of models are trained on 4 NVIDIA A4000 GPUs. All hyperparameters are listed in Table 5. For each classifier, we use domain-specific datasets as positive samples and non-medical datasets as negative samples. For the binary medical/non-medical classifier used in Section 8.2, we use all medical domain datasets as positive samples, and the same group of non-medical datasets as negative samples. See Table 10 for an overview of the datasets used to train these classifiers.
Domain Dataset Total samples Train Test Image Size
CT LIDC-IDRI [11] 10005 7004 3002 512 $\times$ 512
TCGA-LUAD [5] 48931 34252 14679 512 $\times$ 512
WORD [80] 30495 21347 9149 512 $\times$ 512
Positive Videos 1612 1128 484
X-ray ChestX-ray14 [130] 112120 78484 33636 1024 $\times$ 1024
GRAZPEDWRI-DX [88] 20327 14229 6098 660 $\times$ 1660
Shoulder X-ray Classification [23] 841 589 252
Digital Knee X-ray [43] 1650 1155 495 300 $\times$ 162
MURA [105] 40561 28393 12168 1500 $\times$ 2000
Positive Videos 692 484 208 1440 $\times$ 1080
MRI fastMRI [137] 58847 41193 17654 320 $\times$ 320
Duke-Breast-Cancer-MRI [110] 922 645 277 256 $\times$ 256
Medical Segmentation Decathlon [8] 2633 1843 790 256 $\times$ 256
Positive Videos 118 83 35
Dermatology Dermnet [41] 19500 13650 5850 720 $\times$ 472
DDI [31] 656 459 197 300 $\times$ 300
7-point [65] 2045 1432 614 480 $\times$ 720
ISIC [108] 33126 23188 9938
Fitzpatrick 17k [44] 16577 11604 4973
HAM10000 [123] 10015 7011 3005 800 $\times$ 600
Endoscopy KVASIR [99] 8000 5600 2400 720 $\times$ 576
ITEC LapGyn4 [73] 59439 41607 17832 256 $\times$ 256
Red Lesion Endoscopy [29] 3895 2727 1169 320 $\times$ 320
FetReg [15] 12334 8634 3700
TMEDOM [7] 956 669 287
Positive Videos 9496 6647 2849
US COVID-19 Ultrasound [19] 59 41 18
BUSI [4] 780 546 234 500 $\times$ 500
DDTI [97] 134 94 40 560 $\times$ 360
MMOTU [142] 1639 1147 492 3308̃88 $\times$ 2186̃57
HC18 [125] 1334 934 400
EchoNet-Dynamic [93] 10030 7021 3009 112 $\times$ 112
Positive Videos 1874 1312 562
Dentistry Panoramic radiography [106] 598 419 179 2041 $\times$ 1024
ODSI-DB [53] 316 221 95
DENTEX 2023 [47] 2332 1632 700
Dental Calculus [98] 220 154 66
Vident-lab [64] 15110 10577 4533 416 $\times$ 320
Dental condition [111] 1296 907 389 612 $\times$ 408
Oral cancer [135] 144 101 43
Dental cavity [112] 176 123 53
Surg SARAS-ESAD [16] 27175 19023 8153 1920 $\times$ 1080
CholecSeg8k [50] 8080 5656 2424 854 $\times$ 480
DeSmoke-LAP [95] 6000 4200 1800
Surgical Hands [78] 2838 1987 851
m2caiSeg [82] 307 215 92 716 $\times$ 402
NeuroSurgicalTools [20] 2476 1733 743 612 $\times$ 460
ROBUST-MIS 2019 [107] 10000 7000 3000 960 $\times$ 540
Optha Cataracts [6] 35127 24589 10538 1920 $\times$ 1080
Ocular Disease Recognition [2] 3358 2351 1007 512 $\times$ 512
MeDAL Retina [89] 2181 1527 654 768 $\times$ 768
RFMID [94] 3200 2240 960 2144 $\times$ 1424
Glaucoma Detection [1] 650 455 195 3072 $\times$ 2048
DRIVE [121] 40 28 12 584 $\times$ 565
Mammo CBIS-DDSM [114] 10239 7167 3072
CDD-CESM [67] 2006 1404 602 2355x1315
CMMD [30] 5202 3641 1561
Genmed LAION [116] 10861 7603 3258
Non-medical Celeb [77] 202599 60780 28364 178 $\times$ 218
Places [143] 10624928 637496 399497 200 $\times$ 200
AI2D [66] 4903 3432 1471
DocFig [58] 33028 26422 6606
SciFig-Pilot [62] 263952 211162 52790
SlideImages [87] 3452 2762 690
TextVQA [120] 25119 20095 5024
SlideShare-1M [9] 977605 782084 195521
Negative Videos 23956 19165 4791
EgoHands [14] 4800 3840 960 720 $\times$ 1080
11k Hands [3] 11076 8861 2215 1600 $\times$ 1200
IPN Hand [17] 95021 76017 19004 640 $\times$ 480
Table 10: Datasets used to train ResNet50 and ViT-Small medical image classifiers, used in Section 7.2 and Section 8.2.